# Multiplier Between Generalized Toeplitz Kernels

### v1

@@ Abstract.
We develop a structural classification of multipliers between generalized Toeplitz kernels, extending work of Fricain and Rishika. Our results establish new equivalences between multiplier algebras and Carleson-type embeddings, linking them to Beurling Malliavin densities, Pólya sequences, and the spectral theory of entire functions.
@@ 1. Introduction
A function $w\in H^{2}$ induces a Carleson measure for for $\ker T_{g}$ if $\ker T_{g}\subseteq L^{2}(wdm)$. Identifying functions that serve as Carleson measures for Toeplitz kernels remains analytically intricate due to the complex internal structure of these spaces. Consequently, characterizations involving such measures often prove difficult to apply in practice. To address this issue, Fricain and Rishika [9] introduced a simplified framework for characterizing multipliers between model spaces-particularly when the underlying inner functions $U$ and $V$ are meromorphic. Their approach avoids the use of Carleson measures entirely under certain structural assumptions, as captured in the following result:
@@ A Blaschke product is given by
where $\{a_{n}\}\subset\mathbb{D}$ is a sequence satisfying the Blaschke condition
@@ 2.1. Toeplitz Operators and Generalized Kernels
where $P_{+}$ denotes the orthogonal projection of $L^{2}(\mathbb{T})$ onto $H^{2}$. The kernel $\ker T_{\phi}$ encodes deep information about the phase function of $\phi$ and is intimately connected to problems in uniqueness, spectral gaps, and rigidity. In particular, kernels of Toeplitz operators can be viewed as reproducing kernel Hilbert spaces whose structure reflects the fine geometry of singular measures and zero configurations.
To generalize this framework, let $E_{1}\subset L^{2}(\mathbb{T})$ be a closed subspace and $E_{2}$ a simply invariant subspace. We define the generalized Toeplitz operator
$T_{\phi}^{E_{1},E_{2}}:E_{1}\to E_{2},\qquad T_{\phi}^{E_{1},E_{2}}:=P_{E_{2}}(M_{\phi}|_{E_{1}}),$
where $M_{\phi}$ is multiplication by $\phi$ and $P_{E_{2}}$ the orthogonal projection onto $E_{2}$. Such generalized operators arise naturally in the context of Wiener-Hopf theory, and in the study of non-selfadjoint operator algebras.
A natural question arises from the earlier definition of Toeplitz operators: which functions $\varphi\in L^{\infty}(\mathbb{T})$ ensure that $\varphi H^{2}\subset H^{2}$? This leads to the following definition: if $X$ is a Banach space of analytic functions on the unit disk $\mathbb{D}$, an analytic function $\varphi$ on $\mathbb{D}$ is considered a multiplier for $X$ if $\varphi X\subset X$. The space of multipliers of the space $X$ is denoted by $\mathcal{M}(X)$, and it has been established that $\mathcal{M}(H^{2})=H^{\infty}$ [12, 10]. Over the past four decades, model spaces have been extensively studied, revealing deep connections to various domains within complex analysis and operator theory. Consequently, some researchers have also explored the multipliers of model spaces. While the multipliers of $K_{u}$ may not possess the same immediate allure, the investigation of analytic functions $\phi$ on $\mathbb{D}$ that map one model space $K_{u}$ to another model space $K_{v}$ remains a compelling and rich area of inquiry [9, 3, 1].
@@ Acknowledgement
We would like to thanks to Dr. A. K. Bhardwaj for his valuable suggestions and remarks.

### v2

@@ Abstract.
We develop a structural classification of multipliers between generalized Toeplitz kernels, extending the work of Fricain and Rupam. Our results establish new equivalences between multiplier space and Carleson-type embeddings, linking them to Beurling Malliavin densities, Pólya sequences, and the spectral theory of entire functions.
@@ 1. Introduction
A function $w\in H^{2}$ induces a Carleson measure for for $\ker T_{g}$ if $\ker T_{g}\subseteq L^{2}(wdm)$. Identifying functions that serve as Carleson measures for Toeplitz kernels remains analytically intricate due to the complex internal structure of these spaces. Consequently, characterizations involving such measures often prove difficult to apply in practice. To address this issue, Fricain and Rupam [9] introduced a simplified framework for characterizing multipliers between model spaces-particularly when the underlying inner functions $U$ and $V$ are meromorphic. Their approach avoids the use of Carleson measures entirely under certain structural assumptions, as captured in the following result:
@@ A Blaschke product is given by
where $\{a_{n}\}\subset\mathbb{D}$, with $a_{n}\neq 0$ is a sequence satisfying the Blaschke condition
@@ 2.1. Toeplitz Operators and Generalized Kernels
where $P_{+}$ denotes the orthogonal projection of $L^{2}(\mathbb{T})$ onto $H^{2}$. The kernel $\ker T_{\phi}$ encodes deep information about the phase function of $\phi$ and is intimately connected to problems in uniqueness, spectral gaps, and rigidity.
A natural question arises from the earlier definition of Toeplitz operators: which functions $\varphi\in L^{\infty}(\mathbb{T})$ ensure that $\varphi H^{2}\subset H^{2}$? This leads to the following definition: if $X$ is a Banach space of analytic functions on the unit disk $\mathbb{D}$, an analytic function $\varphi$ on $\mathbb{D}$ is considered a multiplier for $X$ if $\varphi X\subset X$. The space of multipliers of the space $X$ is denoted by $\mathcal{M}(X)$, and it has been established that $\mathcal{M}(H^{2})=H^{\infty}$ [12, 10]. Over the past four decades, model spaces have been extensively studied, revealing deep connections to various domains within complex analysis and operator theory. Consequently, some researchers have also explored the multipliers of model spaces. While the multipliers of $K_{u}$ may not possess the same immediate allure, the investigation of analytic functions $\phi$ on $\mathbb{D}$ such that $M_{\phi}$ map one model space $K_{u}$ to another model space $K_{v}$ remains a compelling and rich area of inquiry [9, 3, 1].
@@ Acknowledgement
We would like to thank Dr. A. K. Bhardwaj for his valuable suggestions and insightful remarks.